### v1

@@ 2 Related Works
Online Reinforcement Learning, where an agent continuously learns from ongoing interaction with the environment, instead of using memory replay, has been successfully applied to various robotic tasks. Zhang, Pang, and Hu [18] presented a novel approach to solving the trajectory tracking problem for manipulators and mobile robots using PPO with Generalized Advantage Estimation and Long Short-Term Memory networks as actor and critic. Similarly, Lopes et al. [19] explored the application of Proximal Policy Optimization to control a quadrotor, demonstrating that model-free online reinforcement learning can effectively train a reliable control policy for quadrotor position control. In the humanoid robots area, Melo et al. [20] successfully utilized PPO to develop running skills in a humanoid robot, using the RoboCup 3D Soccer Simulation environment, where humanoid robots compete in simulated soccer matches. Expanding the use of online reinforcement learning to other areas, Proctor [21] used the algorithm for localizing and searching for lost nuclear sources in a designated area, which is vital for societal safety and reducing human harm.
To the best of our knowledge, no efforts have been made to apply KANs as alternatives to MLPs in reinforcement learning, making this study a pivotal contribution.
@@ 3.1 Kolmogorov-Arnold Networks
Comparatively, KANs and MLPs differ as follows:
  • KANs feature learnable activation functions on edges, while MLPs have fixed activation functions on nodes.
  • KANs use splines to represent weights, enhancing their capacity to approximate complex functions with fewer parameters.
  • Both KANs and MLPs can be extended to multiple layers.
@@ 4.1 Network Architectures
For the Kolmogorov-Arnold Network, we set the parameters $k=2$ and $g=3$ based on these preliminary findings, ensuring the network’s ability to approximate the continuous functions involved in control tasks while maintaining low computational complexity. The number of parameters in the KAN configurations is significantly lower compared to the MLP counterparts, highlighting the efficiency of KANs. The average number of parameters for each configuration across the environments is shown in Table 1.
@@ 5.2 Discussion
The results from our experiments indicate that KANs, when integrated into the PPO framework, can achieve performance comparable to traditional MLPs with significantly fewer parameters. This finding is particularly noteworthy as it demonstrates the efficiency of KANs in reducing the computational complexity of reinforcement learning models without compromising on performance.
@@
While KAN-based models generally matched or exceeded the performance of MLP-based models, it is important to acknowledge the contexts in which these improvements were observed. For instance, in tasks like HalfCheetah-v4 and Hopper-v4, the KAN-based PPO not only achieved higher rewards but also maintained a lower parameter count, highlighting the efficiency of KANs. However, in some environments like Pusher-v4 and Walker2d-v4, the performance improvements were less pronounced, or even worse, indicating that the benefits of KANs may vary depending on the specific task and its complexity. This suggests that maybe adjusting parameters, using a different grid search, or incorporating more neurons might yield better performance in these specific cases.
@@
The lower parameter count in KANs translates to reduced memory usage and potentially faster training times, making them advantageous for deployment in resource-constrained environments. For example, the KAN (k=2, g=3) architecture required significantly fewer parameters compared to the MLP (a=2, c=2) configuration, averaging 348 parameters versus 5348, respectively. This substantial reduction in parameter count underscores the efficiency of KANs in achieving comparable performance with fewer resources.
@@
An additional aspect to consider is the prune operation depicted in Figure 3, which involves reducing the number of parameters in the network by eliminating less important ones. Pruning can be particularly beneficial for KANs due to their already reduced parameter count, allowing for even more streamlined models that maintain high performance. The KAN architecture of Swimmer-v4, as shown in the figure, illustrates this pruning process, optimizing the network by focusing on the most critical parameters. Additionally, the inherent interpretability of KANs, based on the Kolmogorov-Arnold representation theorem, provides significant benefits for understanding and debugging the learned policies. This interpretability is crucial in applications where model transparency and explainability are important. However, it is essential to continue exploring and optimizing KAN architectures to fully harness their potential in various reinforcement learning applications.
@@
Despite the efficiency and interpretability benefits, KANs currently face limitations in computational speed. In our evaluation of 1000 steps on the HalfCheetah environment, we observed that the MLP with 2 hidden layers of 64 units completed the task significantly faster, with an average elapsed time of 0.37 seconds, compared to the KAN with no hidden layers (k=2, g=3), which took an average of 3.39 seconds. This increased computational complexity arises from the use of B-splines as activation functions, which require more intricate calculations. The MLP’s architecture, while parameter-heavy, benefits from highly optimized implementations in libraries such as Torch, resulting in faster execution times. As KANs are relatively novel, there is a lack of specialized optimization techniques and frameworks designed specifically for them. This discrepancy highlights the potential need for further research and development to enhance the computational efficiency of KANs, making them more viable for real-time applications.
Figure 3: KAN architecture of Swimmer-v4, illustrating where the pruning is applied in order to enhance network efficiency.

### v3

@@ 2 Related Works
Online Reinforcement Learning, where an agent continuously learns from ongoing interaction with the environment, instead of using memory replay, has been successfully applied to various robotic tasks. In the humanoid robots area, Melo et al. [18] successfully utilized PPO to develop running skills in a humanoid robot, using the RoboCup 3D Soccer Simulation environment, where humanoid robots compete in simulated soccer matches. Expanding the use of online reinforcement learning to other areas, Proctor [19] used the algorithm for localizing and searching for lost nuclear sources in a designated area, which is vital for societal safety and reducing human harm.
To the best of our knowledge, no efforts have been made to apply KANs as alternatives to MLPs in online reinforcement learning, making this study a pivotal contribution.
@@ 3.1 Kolmogorov-Arnold Networks
3.1.1 Grid and Order in KANs
In KANs, the spline functions $\varphi$ are defined over a discretized domain, commonly referred to as a grid. The grid determines the points at which the spline functions are evaluated, essentially specifying the resolution of the function approximation. The order parameter controls the degree of the B-splines used to represent the functions. For example, an order of 1 corresponds to linear splines, while higher orders correspond to splines with more complex shapes, allowing for more flexible function approximations.
The combination of the grid and order parameters defines the complexity and capacity of the KANs to model intricate functional relationships. A finer grid with higher-order splines enables more accurate approximations of complex functions, at the cost of increased computational complexity.
3.1.2 Comparison of KANs and MLPs
  • Learnable Activation Functions: KANs feature learnable activation functions on edges, while MLPs have fixed activation functions on nodes.
  • Spline-Based Weights: KANs use splines to represent weights, enhancing their capacity to approximate complex functions with fewer parameters.
  • Layer Extension: Both KANs and MLPs can be extended to multiple layers.
@@ 4.1 Network Architectures
Figure 2: KAN architecture of the environment DMC Proprio Swimmer-v4.
For the Kolmogorov-Arnold Network, we set the parameters $k=2$ and $g=3$ based on these preliminary findings, ensuring the network’s ability to approximate the continuous functions involved in control tasks while maintaining low computational complexity. The number of parameters in the KAN configurations is significantly lower compared to the MLP counterparts, highlighting the efficiency of KANs. The average number of parameters for each configuration across the environments is shown in Table 1. An Kolmogorov-Arnold Network example is depicted in Fig.2.
@@ 5.2 Discussion
The results from our experiments indicate that KANs, when integrated into the PPO framework, can achieve performance comparable to traditional MLPs with significantly fewer parameters. This demonstrates the potential efficiency of KANs in reducing the computational complexity of reinforcement learning models without compromising performance.
@@
While KAN-based models generally matched or exceeded the performance of MLP-based models, these improvements were context-dependent. For example, in tasks such as HalfCheetah-v4 and Hopper-v4, the KAN-based PPO not only achieved higher rewards but also maintained a lower parameter count, highlighting its efficiency. However, in environments like Pusher-v4 and Walker2d-v4, the performance gains were less pronounced or even inferior, suggesting that the benefits of KANs can vary depending on task complexity. Adjustments in parameters, grid searches, or incorporating additional neurons could potentially enhance performance in these cases.
@@
The reduced parameter count in KANs translates to lower memory usage and potentially faster training times, making them suitable for deployment in resource-constrained environments. For instance, the KAN (k=2, g=3) architecture required significantly fewer parameters compared to the MLP (a=2, c=2) configuration, averaging 348 versus 5348 parameters, respectively. This substantial reduction underscores the efficiency of KANs in achieving comparable performance with fewer resources.
@@
Additionally, pruning operations—where less important parameters are eliminated—can further streamline KAN models, enhancing efficiency without compromising performance. This pruning process is particularly beneficial for KANs due to their already reduced parameter count, allowing the model to focus on the most critical parameters. The inherent interpretability of KANs, based on the Kolmogorov-Arnold representation theorem, also provides advantages for understanding and debugging learned policies, which is crucial in applications where model transparency and explainability are important. However, continued exploration and optimization of KAN architectures are necessary to fully leverage their potential across various reinforcement learning applications.
@@
KANs, while providing significant benefits in terms of parameter count and interpretability, currently face challenges in computational speed. In our evaluation of 1000 steps on the HalfCheetah environment, the MLP with 2 hidden layers of 64 units completed the task significantly faster, averaging 0.37 seconds compared to the KAN (k=2, g=3), which took 3.39 seconds. This slower performance stems from the complex calculations required by the B-splines used as activation functions, which, despite reducing the number of parameters, introduce a trade-off between parameter efficiency and execution speed—especially critical in real-time applications. MLPs, though more parameter-heavy, benefit from highly optimized implementations in libraries like Torch, resulting in faster execution times. The novelty of KANs means there is a lack of specialized optimization techniques, presenting a clear need for future research to enhance their computational efficiency.
Addressing these computational bottlenecks will be essential to fully leverage the potential of KANs in reinforcement learning tasks. Optimizing spline calculations and developing KAN-specific computational strategies could improve their competitiveness in terms of execution speed, making them more viable for applications that demand real-time decision-making while maintaining their advantages in memory usage and training efficiency. Future work should focus on overcoming these challenges to realize the full capabilities of KANs in practical settings.